# Aligning Machine and Human Visual Representations across Abstraction Levels

### v2

@@ 4.4 Alignment with conceptual hierarchy
All data—the Levels dataset and the synthetically created AligNet data—will be made publicly available on GIN upon publication. Prior to publication, the Levels dataset is available upon request (please contact the corresponding authors) and will soon be made available to the public. Both the analysis and the experiment code for Levels are hosted on GitHub and archived on the Zenodo platform.
@@ A.1.1 Representational alignment
We train a VICE model on the official THINGS train triplet dataset using the (default) hyperparameters recommended by the authors. To capture the uncertainties in human triplet responses, VICE learns a mean, $\mu\in\mathbb{R}^{m\times d}$, and a variance, $\sigma\in\mathbb{R}^{m\times d}$, for each object image and each object dimension respectively. Therefore, the set of VICE parameters is defined as $\theta=\{\mu,\sigma\}$. VICE uses the reparameterization trick [Kingma and Welling, 2013; Rezende et al., 2014] to generate an embedding matrix ${\bm{Y}}\in\mathbb{R}^{m\times d}$, ${\bm{Y}}_{\theta,\epsilon}=\mu+\sigma\odot\epsilon$, where $\epsilon\in\mathbb{R}^{m\times d}$ is entrywise $\mathcal{N}(0,1)$, and $\odot$ denotes the Hadamard (element-wise) product.
After convergence, we can use a VICE model to obtain a posterior probability distribution for each triplet in the data. We approximate the probability distribution using a Monte Carlo estimate [Graves, 2011; Blundell et al., 2015; Blei et al., 2016] from $R$ samples ${\bm{Y}}^{(r)}={\bm{Y}}_{\hat{\theta},\epsilon^{(r)}}$ for $r=1,\ldots,R$, yielding
@@ A.1.2 Data generation
Image data. For creating AligNet, we use the publicly available ImageNet database [Deng et al., 2009]. ImagNet is a natural image dataset with $\approx 10^{6}$ training data points and $1000$ image categories [Russakovsky et al., 2014]. The categories are almost equally distributed in the data with small variations in the number of images between the different classes. Hence, ImageNet can be considered a highly balanced dataset. ImageNet has been the dominant image dataset for training large computer vision models until the advent of image/text multi-modal training a few years ago. Although to date larger image datasets exists, ImageNet is still is one of the largest open-source and most widely used image datasets in the field of Computer Vision.
@@
We cluster the representation spaces ${\bm{Z}}$ and ${\bm{Z}}^{\prime}\coloneqq\left({\bm{W}}{\bm{Z}}^{\top}+[{\bm{b}}_{1},\ldots,{\bm{b}}_{m^{\prime}}]\right)^{\top}$ [^2] Each of the $m$ column vectors on the right-hand side is the same ${\bm{b}}$. respectively (see Sec. 3 for how the transformation variables ${\bm{W}}$ and ${\bm{b}}$ are computed) into $c$ representation clusters where $c$ can be regarded as similar to $C$, the number of labels in the original dataset. We use the Elbow criterion to select $c$. For all of our main experiments we set $c=500$. Hence, the ImageNet dataset is transformed into a dataset of image-cluster instead of image-label pairs. Let $\mathcal{D}_{\mathrm{image}}^{\prime}\coloneqq\left(x_{i},y^{\prime}_{i}\right)_{i=1}^{m^{\prime}}$ be the modified ImageNet dataset of image and cluster pairs. After the clustering, we apply the same sampling method as for class-boundary triplets: for each triplet we choose uniformly at random two images without replacement from one cluster and one image from a different cluster. Thus, a triplet generated via cluster-boundary sampling can be defined as the following set $\mathcal{S}\coloneqq\{x_{i},x_{j},x_{k}\}$ with the constraint $\left(y^{\prime}_{i}=y^{\prime}_{j}\neq y^{\prime}_{k}\right)\lor\left(y^{\prime}_{i}\neq y^{\prime}_{j}=y^{\prime}_{k}\right)\lor\left(y^{\prime}_{i}=y^{\prime}_{k}\neq y^{\prime}_{j}\right)$ where instead of the original labels we use the cluster labels for partitioning the data.
Triplet response generation
Let $\mathcal{D}_{\mathrm{triplets}}\coloneqq\left(\{x_{i},x_{j},x_{k}\}_{s}\right)_{s=1}^{n^{\prime}}$ be the dataset of sampled ImageNet triplets for which we want to collect responses using transformed model representations. Note that we can sample an arbitrary number of triplets — upper-bounded by the binomial coefficient $\binom{n^{\prime}}{k}$ with $k=3$ — and can thus set $n^{\prime}$ to essentially any natural number. For the experiments that we report in the main text we set $n^{\prime}=10^{7}$ because we found a larger $n^{\prime}$ to not yield any downstream task improvements. For now, we regard our surrogate model as a blackbox model with transformed ImageNet representations ${\bm{Z}}^{\prime}\coloneqq\left({\bm{W}}{\bm{Z}}^{\top}+[{\bm{b}}_{1},\ldots,{\bm{b}}_{m^{\prime}}]\right)^{\top}\in\mathbb{R}^{m^{\prime}\times p}$ where the affine transformation was found via UD optimization (see Eq. 3).
@@ A.1.4 Surrogate teacher model
We use the SigLIP-So400m variant of SigLIP as our teacher model. The So400m variant uses an optimized ViT backbone whose performance is similar to one of the largest ViTs, ViT-g/14 [Alabdulmohsin et al., 2023] while having fewer parameters and thus being smaller. The number of parameters of SoViT-400m/14 lies somewhere between ViT-L/16 and ViT-g/14. The output dimensionality of the image and text encoder representations of SoViT-400m/14 is $p$=$1152$ each. We align the image encoder representations with human odd-one-out choices using the UD optimization outlined in Eq. 6. This allows us to increase the triplet odd-one-out accuracy of SigLIP-So400m from $44.24\%$ to $61.7\%$ (see rightmost column in Tab. 1) which is close to the human-noise ceiling of $66.67\%$ for THINGS [c.f. Hebart et al., 2020] and thus among the best human-aligned models without AligNet fine-tuning [c.f. Muttenthaler et al., 2023a]. Note that this is a relative increase in performance of $39.47\%$. Throughout this manuscript we use the human-aligned version of SigLIP-So400m as the surrogate teacher model for generating human-like similarity judgments and distilling human-like similarity structure into student VFMs. We select a diverse and representative set of student VFMs.
@@ A.1.5 Student Models
Since previous research has demonstrated that a model’s architecture has no significant impact on the degree of alignment with human similarity judgments [Muttenthaler et al., 2023a, b], we use the same architecture for all student models that we fine-tune on AligNet. Specifically, we use the Vision Transformer [ViT; Dosovitskiy et al., 2020] for the backbone of each student model. We use the ViT rather than CNN-based model because ViTs have recently emerged as the dominant neural network architecture for computer vision application and VFMs. Every large VFM that is used in practice is based on the ViT [Tschannen et al., 2023; Chen et al., 2023; Zhai et al., 2023; Beyer et al., 2024]. Unless otherwise mentioned, we use the base model size, i.e., ViT-B. ViT-B has $12$ attention layers and an internal (hidden) representation size of $p=768$. It has been shown that both the training data and the objective function have a substantial impact on the degree of alignment with human behavior. Thus, we use student models that were trained on different pretraining task with different training data and objective functions.
Supervised pre-training is still the prevailing mode of training computer vision models. Therefore, we trained ViT-B on the popular ImageNet dataset consisting of 1.4M natural images [Deng et al., 2009]. To examine how model performance changes as a function of the size of a model, we train ViT instances of three different sizes on ImageNet: ViT-S/16, ViT-B/16, ViT-L/16. The image patch size is the same for each of those models. To evaluate the effect of AligNet on self-supervised pretraining, we use pretrained DINOv1 [Caron et al., 2021] and DINOv2 [Oquab et al., 2023] models of which DINOv1 was pretrained on ImageNet and DINOv2 was pretrained on a different, larger image dataset as denoted below. In addition, we investigate multimodal training of vision models that add textual information both in the form of image captioning via the CapPa model [Tschannen et al., 2023], as well as contrastive language-image pretraining (CLIP) via SigLIP [Zhai et al., 2023]. The latter model is considered state-of-the-art on many downstream computer vision applications and is used as the image embedding model in modern large visual-language models [VLMs; Chen et al., 2023; Beyer et al., 2024]. The full list of student models that we consider in our analyses is the following:
@@
    • Objective: Self-supervised teacher-student distillation [Oquab et al., 2023]
@@ Appendix B Additional/Detailed Results
In addition to the results that we presented in the main paper, here, we report the same set of results in more detail and add further analyses that we touched upon in the main part but did not go into much detail. We start by presenting findings from a comprehensive set of evaluations that we performed to test the alignment of model representations with human similarity judgments both for similarity judgments collected in previous efforts [c.f. Peterson et al., 2018; Cichy et al., 2019; King et al., 2019; Hebart et al., 2020] and for our own similarity judgment dataset Levels (see Section A.3). Subsequently, we demonstrate benefits of our method for machine learning downstream tasks that test generalization, such as few-shot learning and out-of-distribution detection. Finally, we perform a rigorous qualitative analysis of the changes in the model representations after applying our alignment framework to the various student models.
@@ B.1.1 Detailed results for RSA
In the bottom two rows of Fig. 6 and Fig. 7 respectively we show model performances individually for the single category pairwise similarity rating datasets. We observe that AligNet fine-tuning significantly improved the degree of alignment for all models for the various, animals, and automobiles categories but there does not appear to be a difference in performance between the different model transformation settings for the categories fruits, furniture, and vegetables. We hypothesize that this difference in alignment benefits stems from the fact that animal- and automobile-related concepts are more frequently represented in the THINGS dataset than fruits-, furniture-, and vegetable-related concepts [c.f. Hebart et al., 2019, 2020]. Moreover, THINGS is a human similarity judgment dataset that reflects global coarse-grained semantic, similar to the coarse-grained setting in Levels. Thus, it is rather surprising to find benefits for human local fine-grained semantic by either UD or fine-tuning on AligNet.
@@ B.2.1 Few-shot learning
Few-shot learning Wang et al. [2021] is a common way to evaluate how easily a neural network representation can generalize to new tasks: The goal here is to learn a new task/dataset from very few labelled examples (typically $\leq$10 examples per labelled class), and then evaluate its performance on the full test set of that task. Good few-shot performance is indicative of models that have a solid understanding of many varying concepts, and that their representation is able to quickly integrate/expand to new or related concepts.
We run few-shot learning evaluations on eight different supervised tasks. ImageNet [Russakovsky et al., 2014], Places365 [Zhou et al., 2018] and Cifar100 [Krizhevsky, 2009] test for broad understanding of a wide variety of visual concepts, and gauge how easily a representation can adapt to a new learning task on concepts a (pretrained) model has likely encountered before. We also test datasets that instead require that the model is able to differentiate between very fine-grained differences in related concepts. Concretely, we test datasets such as Caltech-UCSD Birds 200 (CUB-200) Birds [Welinder et al., 2010], Stanford Cars [Krause et al., 2013], Flowers [Nilsback and Zisserman, 2008], Oxford-IIIT Pets [Parkhi et al., 2012] and the UC Merced Land Use Dataset Yang and Newsam [2010]. For our tests, we freeze the model, and train a linear classifier on top of the extracted image representation Alain and Bengio [2016]. On all datasets, we test two few-shot settings: the extreme case where only a single example per class is provided (‘1-shot‘) and the more conservative, but still challenging ‘10-shot‘ challenge.
@@ B.2.3 Ablations
We performed several ablations to see how the AligNet procedure changes if we modify some of its crucial parameters.
Model Capacity
To investigate if AligNet finetuning affects models of different capacity the same way, we compared the performance on the ViT-B model (87M parameters) with a smaller ViT-S (22M parameters) and a larger ViT-L (305M parameters) model. The results in Fig. 12 and Fig. 13 indicate that the positive impact of AligNet increases with model scale. For example, the average 10-shot performance improved by 2.4% points for ViT-S, by 6.3% points for ViT-B and by 8.3% points for ViT-L. Similar results hold up for all datasets we investigated. Our hypothesis is that the larger model capacity afford the model to better absorb the additional high-level information provided by the human similarity judgments distilled into AligNet.
@@
Increasing the Number of Images or Triplets
@@
Using an Unaligned Teacher Model
@@
Investigating different Triplet Sampling Schemes
@@
Fine-tuning directly on THINGS
@@
We have argued that fine-tuning models to better align them with human similarity judgments depends on having sufficiently large datasets, which do not exist yet. This argument has motivated our multi-stage approach of first linearly aligning teacher model and then using this as a surrogate to generate a large corpus of human-like triplets. Yet, if we look at Fig. 14, it seems that decreasing the number of images or the number of triplets by up to two orders of magnitude still retains most of the benefit on downstream performance (in terms of few-shot accuracy). This raises the question, whether our pipeline as a whole might be unnecessary, given that the THINGS [Hebart et al., 2023] dataset already contains 4.70 million human similarity judgments collected via online crowd- sourcing for 1854 object images. In this section, we therefore compare AligNet based training with just using the THINGS odd-one-out triplets directly for fine-tuning of student models. Fig. 17 shows that fine-tuning diretly on THINGS is consistently worse than training on AligNet in terms of ImageNet 1-shot accuracy (x-axis). Importantly, it appears that in almost all cases the performance of the model actually decreased below that of the base model. It is also interesting, that even in terms of THINGS odd-one-out accuracy, the AligNet trained model performs better than the any other model. This is somewhat surprising, because it means that there is an overfitting problem when training on THINGS, that seems to mostly disappear when training on AligNet.

### v3

@@ 4.4 Alignment with conceptual hierarchy
The synthetically created AligNet data will be made publicly available upon publication. We already made the Levels data available on GIN prior to publication: https://doi.org/10.12751/g-node.hg4tdz. Both the analysis and the experiment code for Levels are hosted on GitHub and archived on the Zenodo platform (see https://zenodo.org/doi/10.5281/zenodo.13749101).
@@ A.1.1 Representational alignment
We train a VICE model on the official THINGS train triplet dataset using the (default) hyperparameters recommended by the authors. To capture the uncertainties in human triplet responses, VICE learns a mean, $\mu\in\mathbb{R}^{m\times d}$, and a variance, $\sigma\in\mathbb{R}^{m\times d}$, for each object image and each object dimension respectively. Therefore, the set of VICE parameters is defined as $\theta=\{\mu,\sigma\}$. VICE uses the reparameterization trick [Kingma and Welling, 2014; Rezende et al., 2014] to generate an embedding matrix ${\bm{Y}}\in\mathbb{R}^{m\times d}$, ${\bm{Y}}_{\theta,\epsilon}=\mu+\sigma\odot\epsilon$, where $\epsilon\in\mathbb{R}^{m\times d}$ is entrywise $\mathcal{N}(0,1)$, and $\odot$ denotes the Hadamard (element-wise) product.
After convergence, we can use a VICE model to obtain a posterior probability distribution for each triplet in the data. We approximate the probability distribution using a Monte Carlo estimate [Graves, 2011; Blundell et al., 2015; Blei et al., 2017] from $R$ samples ${\bm{Y}}^{(r)}={\bm{Y}}_{\hat{\theta},\epsilon^{(r)}}$ for $r=1,\ldots,R$, yielding
@@ A.1.2 Data generation
Image data. For creating AligNet, we use the publicly available ImageNet database [Deng et al., 2009]. ImagNet is a natural image dataset with $\approx 10^{6}$ training data points and $1000$ image categories [Russakovsky et al., 2015]. The categories are almost equally distributed in the data with small variations in the number of images between the different classes. Hence, ImageNet can be considered a highly balanced dataset. ImageNet has been the dominant image dataset for training large computer vision models until the advent of image/text multi-modal training a few years ago. Although to date larger image datasets exists, ImageNet is still is one of the largest open-source and most widely used image datasets in the field of Computer Vision.
@@
We cluster the representation spaces ${\bm{Z}}$ and ${\bm{Z}}^{\prime}\coloneqq\left({\bm{W}}{\bm{Z}}^{\top}+\left({\bm{b}}_{1},\ldots,{\bm{b}}_{m^{\prime}}\right)\right)^{\top}$ [^2] Each of the $m$ column vectors on the right-hand side is the same ${\bm{b}}$. respectively (see Sec. 3 for how the transformation variables ${\bm{W}}$ and ${\bm{b}}$ are computed) into $c$ representation clusters where $c$ can be regarded as similar to $C$, the number of labels in the original dataset. We use the Elbow criterion to select $c$. For all of our main experiments we set $c=500$. Hence, the ImageNet dataset is transformed into a dataset of image-cluster instead of image-label pairs. Let $\mathcal{D}_{\mathrm{image}}^{\prime}\coloneqq\left(x_{i},y^{\prime}_{i}\right)_{i=1}^{m^{\prime}}$ be the modified ImageNet dataset of image and cluster pairs. After the clustering, we apply the same sampling method as for class-boundary triplets: for each triplet we choose uniformly at random two images without replacement from one cluster and one image from a different cluster. Thus, a triplet generated via cluster-boundary sampling can be defined as the following set $\mathcal{S}\coloneqq\{x_{i},x_{j},x_{k}\}$ with the constraint $\left(y^{\prime}_{i}=y^{\prime}_{j}\neq y^{\prime}_{k}\right)\lor\left(y^{\prime}_{i}\neq y^{\prime}_{j}=y^{\prime}_{k}\right)\lor\left(y^{\prime}_{i}=y^{\prime}_{k}\neq y^{\prime}_{j}\right)$ where instead of the original labels we use the cluster labels for partitioning the data.
Triplet response generation.
Let $\mathcal{D}_{\mathrm{triplets}}\coloneqq\left(\{x_{i},x_{j},x_{k}\}_{s}\right)_{s=1}^{n^{\prime}}$ be the dataset of sampled ImageNet triplets for which we want to collect responses using transformed model representations. Note that we can sample an arbitrary number of triplets — upper-bounded by the binomial coefficient $\binom{n^{\prime}}{k}$ with $k=3$ — and can thus set $n^{\prime}$ to essentially any natural number. For the experiments that we report in the main text we set $n^{\prime}=10^{7}$ because we found a larger $n^{\prime}$ to not yield any downstream task improvements. For now, we regard our surrogate model as a blackbox model with transformed ImageNet representations ${\bm{Z}}^{\prime}\coloneqq\left({\bm{W}}{\bm{Z}}^{\top}+\left({\bm{b}}_{1},\ldots,{\bm{b}}_{m^{\prime}}\right)\right)^{\top}\in\mathbb{R}^{m^{\prime}\times p}$ where the affine transformation was found via UD optimization (see Eq. 3).
@@ A.1.4 Surrogate teacher model
We use the SigLIP-So400m variant of SigLIP as our teacher model. The So400m variant uses an optimized ViT backbone whose performance is similar to one of the largest ViTs, ViT-g/14 [Alabdulmohsin et al., 2023] while having fewer parameters and thus being smaller. The number of parameters of SoViT-400m/14 lies somewhere between ViT-L/16 and ViT-g/14. The output dimensionality of the image and text encoder representations of SoViT-400m/14 is $p$=$1152$ each. We align the image encoder representations with human odd-one-out choices using the UD optimization outlined in Eq. 6. This allows us to increase the triplet odd-one-out accuracy of SigLIP-So400m from $44.24\%$ to $61.7\%$ (see rightmost column in Tab. 1) which is close to the human-noise ceiling of $66.67\%$ for THINGS [cf. Hebart et al., 2020] and thus among the best human-aligned models without AligNet fine-tuning [cf. Muttenthaler et al., 2023a]. Note that this is a relative increase in performance of $39.47\%$. Throughout this manuscript we use the human-aligned version of SigLIP-So400m as the surrogate teacher model for generating human-like similarity judgments and distilling human-like similarity structure into student VFMs. We select a diverse and representative set of student VFMs.
@@ A.1.5 Student Models
Since previous research has demonstrated that a model’s architecture has no significant impact on the degree of alignment with human similarity judgments [Muttenthaler et al., 2023a, b], we use the same architecture for all student models that we fine-tune on AligNet. Specifically, we use the Vision Transformer [ViT; Dosovitskiy et al., 2021] for the backbone of each student model. We use the ViT rather than CNN-based model because ViTs have recently emerged as the dominant neural network architecture for computer vision application and VFMs. Every large VFM that is used in practice is based on the ViT [Tschannen et al., 2023; Chen et al., 2023; Zhai et al., 2023; Beyer et al., 2024]. Unless otherwise mentioned, we use the base model size, i.e., ViT-B. ViT-B has $12$ attention layers and an internal (hidden) representation size of $p=768$. It has been shown that both the training data and the objective function have a substantial impact on the degree of alignment with human behavior. Thus, we use student models that were trained on different pretraining task with different training data and objective functions.
Supervised pre-training is still the prevailing mode of training computer vision models. Therefore, we trained ViT-B on the popular ImageNet dataset consisting of 1.4M natural images [Deng et al., 2009]. To examine how model performance changes as a function of the size of a model, we train ViT instances of three different sizes on ImageNet: ViT-S/16, ViT-B/16, ViT-L/16. The image patch size is the same for each of those models. To evaluate the effect of AligNet on self-supervised pretraining, we use pretrained DINOv1 [Caron et al., 2021] and DINOv2 [Oquab et al., 2024] models of which DINOv1 was pretrained on ImageNet and DINOv2 was pretrained on a different, larger image dataset as denoted below. In addition, we investigate multimodal training of vision models that add textual information both in the form of image captioning via the CapPa model [Tschannen et al., 2023], as well as contrastive language-image pretraining (CLIP) via SigLIP [Zhai et al., 2023]. The latter model is considered state-of-the-art on many downstream computer vision applications and is used as the image embedding model in modern large visual-language models [VLMs; Chen et al., 2023; Beyer et al., 2024]. The full list of student models that we consider in our analyses is the following:
@@
    • Objective: Self-supervised teacher-student distillation [Oquab et al., 2024]
@@ Appendix B Additional/Detailed Results
In addition to the results that we presented in the main paper, here, we report the same set of results in more detail and add further analyses that we touched upon in the main part but did not go into much detail. We start by presenting findings from a comprehensive set of evaluations that we performed to test the alignment of model representations with human similarity judgments both for similarity judgments collected in previous efforts [cf. Peterson et al., 2018; Cichy et al., 2019; King et al., 2019; Hebart et al., 2020] and for our own similarity judgment dataset Levels (see Section A.3). Subsequently, we demonstrate benefits of our method for machine learning downstream tasks that test generalization, such as few-shot learning and out-of-distribution detection. Finally, we perform a rigorous qualitative analysis of the changes in the model representations after applying our alignment framework to the various student models.
@@ B.1.1 Detailed results for RSA
In the bottom two rows of Fig. 6 and Fig. 7 respectively we show model performances individually for the single category pairwise similarity rating datasets. We observe that AligNet fine-tuning significantly improved the degree of alignment for all models for the various, animals, and automobiles categories but there does not appear to be a difference in performance between the different model transformation settings for the categories fruits, furniture, and vegetables. We hypothesize that this difference in alignment benefits stems from the fact that animal- and automobile-related concepts are more frequently represented in the THINGS dataset than fruits-, furniture-, and vegetable-related concepts [cf. Hebart et al., 2019, 2020]. Moreover, THINGS is a human similarity judgment dataset that reflects global coarse-grained semantic, similar to the coarse-grained setting in Levels. Thus, it is rather surprising to find benefits for human local fine-grained semantic by either UD or fine-tuning on AligNet.
@@ B.2.1 Few-shot learning
Few-shot learning [Wang et al., 2020] is a common way to evaluate how well neural network representations generalize to new tasks: Here, the goal is to learn a new task/dataset from only a few labelled examples (typically $\leq$10 examples per labelled class are used to train a classifier on top of the pretrained representations), and then evaluate its performance on the full test set for that task. Good few-shot performance is indicative of models that have a solid understanding of many different concepts, and that their representations are able to quickly integrate/expand to new or related concepts.
We run few-shot learning evaluations on eight different supervised learning tasks. ImageNet [Russakovsky et al., 2015], Places365 [Zhou et al., 2018] and Cifar100 [Krizhevsky and Hinton, 2009] test for broad understanding of a wide variety of visual concepts, and gauge how easily a representation can adapt to a new learning task on concepts a (pretrained) model has likely encountered before. We also test datasets that instead require that the model is able to differentiate between very fine-grained differences in related concepts. Concretely, we test datasets such as Caltech-UCSD Birds 200 (CUB-200) Birds [Wah et al., 2011], Stanford Cars [Krause et al., 2013], Flowers [Nilsback and Zisserman, 2008], Oxford-IIIT Pets [Parkhi et al., 2012] and the UC Merced Land Use Dataset [Yang and Newsam, 2010]. For our evaluation, we freeze the model weights, and train a linear classifier on top of the extracted image representation [Alain and Bengio, 2017]. On all datasets, we test two few-shot settings: the extreme case where only a single example per class is provided (‘1-shot‘) and the more conservative, but still challenging ‘10-shot‘ challenge.
@@
Tip-Adapter.
Muttenthaler et al. [2023b] have demonstrated that linearly transforming the pretrained representations of state-of-the-art image/text models into a human-aligned concept space can improve upon methods that were specifically designed for improving the performance of pretrained representations in few-shot learning tasks. One of these methods is Tip-Adapter [Zhang et al., 2022]. Tip-Adapter is designed to yield effective few-shot classifiers for image/text models by linearly combining the output of two modules — a zero-shot classifier and a key-value cache model. UD (see Appx. A.1.1) and gLocal [Muttenthaler et al., 2023b] have highly similar learning objectives: both linearly transform the representations of pretrained vision foundation models into a (global) human-aligned concept space while preserving the models’ nearest neighbor structure. Since we find that UD either performs equally well or better in matching the human similarity judgments and predicting the human RTs for the triplets in the Levels data (see Appx. B.2.3) and soft-aligned models outperform the UD transformation across the different human object similarity tasks that we evaluated (see Appx. B.1 for details), we can be confident that soft-aligned image/text models (i.e., image/text models fine-tuned on AligNet) can improve upon Tip-Adapter. For CIFAR-100, the improvements in zero-shot performance of using gLocal in conjunction with Tip-Adapter compated to Tip-Adapter alone ranged from $6.29\%$ for a CLIP-RN50 up to $18.72\%$ for a CLIP ViT-L/14 (see Appendix D in Muttenthaler et al. [2023b]). For other datasets, the improvements in zero-shot performance were similarly high. One caveat with Tip-Adapter is that its application is restricted to models that have a text encoder module, i.e., image/text models. Soft-alignment, on the other hand, does not suffer from the same constraint and can be applied to any vision model.
@@ B.2.3 Ablations
We performed several ablations to examine how the effects of AligNet change if we modify some of its central parameters.
Model Capacity.
To investigate if AligNet finetuning affects models of different capacity/size in a similar way, we compared the performance of the ViT-B model (87M parameters) with the performances of a smaller ViT-S (22M parameters) and a larger ViT-L (305M parameters) model respectively. The results in Fig. 12 and Fig. 13 indicate that the positive impact of AligNet fine-tuning increases with model scale. For example, the average 10-shot performance improved by $2.4\%$ points for ViT-S, by $6.3\%$ points for ViT-B and by $8.3\%$ points for ViT-L. We observe similar results for all datasets we considered. Thus, we hypothesize that a larger model capacity makes it easier for a model to integrate the additional global coarse-grained information provided by AligNet’s human-like similarity judgments.
@@
Increasing the Number of Images or Triplets.
@@
Using an Unaligned Teacher Model.
@@
Investigating different Triplet Sampling Schemes.
@@
Fine-tuning directly on the THINGS data.
@@
We have stressed that fine-tuning models to better align them with human similarity judgments depends on having sufficiently large datasets, which do not exist yet. This argument has motivated our multi-stage approach of first linearly aligning a teacher model and then using this as a surrogate model to generate a large corpus of human-like triplet odd-one-out choices. Yet, if we look at Fig. 14, it seems that decreasing the number of images or the number of triplets by up to two orders of magnitude still retains most of the benefit on downstream task performance (in terms of few-shot accuracy). This raises the question whether our pipeline as a whole might be unnecessary, given that the THINGS dataset [Hebart et al., 2023] already contains 4.70 million human similarity judgments collected via online crowd- sourcing for 1854 object images.
Therefore, here we compare soft-alignment with using the THINGS odd-one-out triplets directly for fine-tuning student models. Fig. 17 shows that fine-tuning directly on THINGS is consistently worse compared to training on AligNet in terms of ImageNet 1-shot accuracy ($x$-axis). Importantly, it appears that in almost all cases the performance of the model actually decreases below that of the base model. It is further interesting to note that even in terms of THINGS odd-one-out accuracy, the AligNet trained models perform better than the any other model. This is somewhat surprising, because it means that there clearly exists an overfitting problem when training on the THINGS data (even with strong $\ell_{2}$-regularization), which seems to mostly disappear when using AligNet for fine-tuning. One reason for this could be the differences in the sampling procedure (cluster-based vs. random sampling) and the loss function (soft-choices vs. hard-choices) between AligNet and THINGS.
Figure 18: gLocal vs. UD on the Levels data. Here, we compare gLocal against UD for the different abstraction levels in the Levels dataset. A: Fraction of triplets where humans and models selected the same odd-one-out image. B: Spearman rank correlations between the human disagreement levels (using the five responses per triplet) and the model uncertainties over the triplets. C: Spearman rank correlations between the aggregated human response times (in log-space) and the model triplet uncertainties. Dashed horizontal lines in the panels of the top row indicate random guessing.
Uncertainty Distillation vs. gLocal. Here, we compare the Uncertainty Distillation (UD) transformation that we introduced in Appx. A.1.1 against gLocal [cf. Muttenthaler et al., 2023b]. The learning objectives of the two linear transformations are highly similar and the goal is the same: the representations of a pretrained vision foundation model are linearly transformed into a global coarse-grained human object similarity space using the THINGS dataset [Hebart et al., 2020] while the nearest neighbor structure of the model’s representation space is preserved. However, the difference between gLocal and UD is that UD directly injects the human uncertainty estimates for the triplet odd-one-out choices via approximate Bayesian inference [cf. Muttenthaler et al., 2022] into a model’s representation space and, thus, uses soft rather than hard triplet choices (see Appx. A.1.1 for details). We use the Levels datasets (see Appx. A.3) for evaluating the UD and gLocal transformations, comparing the same set of models that were evaluated in Muttenthaler et al. [2023b]. The gLocal transformations for those models are publicly available via the Python library thingsvision [Muttenthaler and Hebart, 2021] which we used for this analysis. In Fig. 18 we see that the UD transformation is either equally accurate or better in aligning the model’s representations with the human similarity judgments, and predicting their disagreement levels and aggregated RTs respectively in the different abstraction level compared to gLocal. As before, human RTs are measured in log-space and model uncertainties are measured as discrete Shannon entropy over the triplet odd-one-out choices. Note that UD is computationally more efficient than gLocal because it does not require an extra set of images—in addition to the THINGS data—to evaluate the preservation of a representation’s nearest neighbor structure which the multi-objective of gLocal necessitates [cf. Muttenthaler et al., 2023b].